ADNAN MUHAMMED

Grade 12,

A frictionless wire AB is fixed on a sphere of radius R. A very small spherical ball slips on this wire. The time taken by this ball to slip from A to B is

A frictionless wire AB is fixed on a sphere of radius R. A very small spherical ball slips on this wire. The time taken by this ball to slip from A to B is

Grade:11

3 Answers

Vikas TU
6 years ago
time taken by this ball to slip from A to B is:
AB = R/(cos theta)          theta = 45o                            
AB = √2 R = 2R/√2 = 2R cosѲ
  • 2R cosѲ = 1/2x gcosѲ x t2  
t = 2 √R/g
sheela pundir
6 years ago
a=gcostheta and from triangle ABC AB=2Rcos thetaH=1/2gt2 2Rcostheta = 1/2gcostheta t2t=2root r/g Hope it helpssssss u.........
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

time taken by this ball to slip from A to B is:
AB = R/(cos theta) theta = 45o
AB = √2 R
= 2R/√2
= 2R cosѲ2R cosѲ
= 1/2x gcosѲ x t2t
= 2 √R/g

Thanks and Regards